How old are you Grand ma???Little boy said
Grand Ma Reply.....
"next year" said the old lady," I will be six times as old as i am now, less six times as old as I was 10 years ago"
How old is GrandMa???

Answer by nerdybill(7384)

How old are you Grand ma???Little boy said
Grand Ma Reply.....
"next year" said the old lady," I will be six times as old as i am now, less six times as old as I was 10 years ago"
How old is GrandMa???
.
Let x = Grandma's age
then
"next year age of Grandma" = "six times as old as i am now" - "six times as old as I was 10 years ago"
.
Gives us our equation:
x+1 = 6x - 6(x-10)
x+1 = 6x - 6x+60
x+1 = 60
x = 59 years old